Circular No. 5649

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GEMINGA The Compton Observatory EGRET Team (H. A. Mayer-Hasselwander,

K. Brazier, G. Kanbach, and C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik; D. L. Bertsch, B. L. Dingus, C. E. Fichtel, R. C. Hartman, S. D. Hunter, P. W. Kwok, J. R. Mattox, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; J. Chiang, J. Fierro, Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; and E. Schneid, Grumman Aerospace Corporation) communicates: “Since deriving the previous EGRET ephemeris (IAUC 5583), we have obtained 4 more weeks of low-sensitivity exposure to Geminga (30^{o} off-axis). This allows an accurate ephemeris to be derived from EGRET data alone: T_{0} = 1991 May 24.0 Barycentric Dynamical Time (TDB), f_{0} = 4.217 674 997 9 Hz, f_{1} = –1.9526(7) × 10^{−13} Hz/s. The phase of peak 1 is 0.64(1) at T_{0} at the barycenter and 0.21(1) on 1991 May 24.0 UT at the geocenter. This ephemeris is expected to be correct to within 0.1 of a revolution for the interval 1990.3-1993.7. The EGRET data indicate that the second frequency derivative is < 2 × 10^{−23} Hz s^{−2}. The frequency derivative f_{1} is inconsistent with the value given in the previous EGRET ephemeris, which was obtained by an extrapolation of the COS-B ephemeris (IAUC 5541). Thus, the phase of the previous EGRET ephemeris is correct to within 0.1 of a revolution for a shorter interval (1990.6-1992.9) than previously specified. The epoch of the previous EGRET ephemeris was specified as UT. However, it was actually TDB. We apologize for the confusion this has caused. The position of Geminga is α = 6^{h}33^{m}54^{s}.02, δ = +17^{o}46′11′′.5 (equinox J2000.0)”

PERIODIC COMET SWIFT-TUTTLE (1992t) Further total visual magnitude estimates (cf. IAUC 5645): Oct. 30.06

UT, 6.2 (B. H. Granslo, Fjellhamar, Norway, 0.06-m refractor); 31.83, 5.9 (S. Garro, Merlette, France, naked eye); Nov. 1.99, 6.0 (B. G. Marsden, Lexington, MA, 7×50 binoculars); 3.40, 5.8 (M. Ohkuma, Mt. Fuji, Japan, 10×70 binoculars; 1^{o} tail in p.a. 40^{o}); 5.09, 5.4 (W. G. Dillon, Missouri City, TX, 7×50 binoculars); 5.78, 5.5 (M. Tombelli, Montelupo, Italy, 11×80 binoculars).

1992 November 6 Daniel W. E. Green